# Displacement techniques in bounded cohomology

### v1

@@ 2.2.4. Conjugates in commuting $\mathbb{Z}$-conjugate
We say that $\Gamma$ has conjugates in commuting $\mathbb{Z}$-conjugate (abbreviated $\operatorname{c.c.\mathbb{Z}\ c.}$) if there exists a subgroup $\Lambda\leq\Gamma$ and an element $t\in\Gamma$ such that:
@@ 3.1. Correctness of Figure 1
Consider the group $\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{2})$. It has commuting $\mathbb{Z}$-conjugates [CFFLM23, Corollary 5.20], as well as commuting $\mathbb{Z}/2$-conjugates [CFFLM23, Corollary 5.12]. It is straightforward to adapt the proof to show that it has commuting $\mathbb{Z}/n$-conjugates for every $n\geq 2$.
@@
We fix a sequence $\textbf{n}=(n_{1},n_{2},\ldots)$, where each $n_{i}\in\mathbb{Z}_{\geq 2}$. Set $\Gamma_{0}$ to be an arbitrary finitely generated non-abelian group, and define by induction $\Gamma_{i+1}\coloneqq\Gamma_{i}\wr\mathbb{Z}/n_{i+1}$. The standard embeddings $\Gamma_{i}\to\Gamma_{i+1}$ make this into a directed system of groups, and we denote by $\Gamma_{\mathbf{n}}$ their directed union.
@@
Let $H_{1}<\operatorname{\textup{Homeo}_{c}}(\mathbb{R})$ be the copy of Thompson’s group $F$ acting on $(0,1)$ and fixing the rest of $\mathbb{R}$ pointwise. Note that $H_{1}$ is finitely generated and acts minimally on $(0,1)$. If $H_{1}$ commutes with $t^{n}$, it follows in particular that $t^{n}(0,1)=(0,1)$, and thus $t^{n}(0)=0,t^{n}(1)=1$. Since $t$ preserves the orientation, we deduce that $t(0)=0$ and $t(1)=1$. Thus ${}^{t}H_{1}$ is also supported on $(0,1)$ and acts minimally on it. It follows that $H_{1}$ and ${}^{t}H_{1}$ do not commute. ∎
@@
Now the proof of Proposition 3.5 shows that there exists a finitely generated subgroup $H_{1}<\Gamma_{1}$, such that if $[H_{1},t_{1}^{n}]=1$ for some $n\geq 2$, then $[H_{1},{}^{t_{1}}H_{1}]\neq 1$. The assumption on $\Gamma_{2}$ implies that there exists a finitely generated subgroup $H_{2}<\Gamma_{2}$ such that for all $t_{2}\in\Gamma_{2}$ there exists $p\geq 1$ such that $[H_{2},{}^{t_{2}^{p}}H_{2}]\neq 1$.
@@ 3.2. More examples with commuting $\mathbb{Z}/2$-conjugates but not commuting $\mathbb{Z}$-conjugates
We will focus on an example of great interest in $K$-theory and homological stability: The infinite general linear group $\operatorname{GL}(\mathbb{R})$. This situation showcases how important the notion of commuting cyclic conjugates is in order to obtain $\operatorname{\mathfrak{X}^{\textup{sep}}}$-bounded acyclicity for classical groups, instead of restricting to the property of commuting $\mathbb{Z}$-conjugates. Our proof is based on the fact that the displacing element in $\operatorname{GL}(\mathbb{R})$ has to conserve the dimension (of the matrices in the finitely generated subgroup). For this reason, a similar argument does also apply to some $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic groups that preserve a quantity [CFFLM23], including e.g.
@@
The fact that $\operatorname{GL}(\mathbb{R})$ has commuting cyclic conjugates has already been proved [CFFLM23, Corollary 5.46], so we are left to show here that $\operatorname{GL}(\mathbb{R})$ does not have commuting $\mathbb{Z}$-conjugates.
@@ 3.3. Correctness of Figure 2
It is easier to work with the finitely generated groups $\Gamma_{i}$ than with the groups $G_{i}$. In this respect, the following lemma is useful:
@@ 3.5. Mitotic groups
Despite the similarity of the problem, adapting the argument from Proposition 3.19 to the case of commuting cyclic conjugates presents a challenge. Namely, a crucial step in the proof that $\beta(\Gamma)$ does not have commuting conjugates is the fact that the image $b^{i-1}(\Gamma)\to b^{i}(\Gamma)$ intersects trivially all edge stabilizers in the corresponding Bass–Serre tree. This is however not the case with $m^{i-1}(\Gamma)\to m^{i}(\Gamma)$, where the image is identified with an edge stabilizer, and in fact can be shown to fix infinitely many edges.
@@ 4. Problems
A natural candidate of a binate group that is not $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic is the binate tower construction, since such groups do not have commuting cyclic conjugates: In fact they do not even have commuting conjugates (Proposition 3.19).
@@
Another possible approach for constructing a boundedly acyclic group which is not $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic goes via commensurability. Indeed, let $\Lambda<\Gamma$ be a finite index subgroup. Then for every Banach $\Gamma$-module $E$, the restriction $H^{n}_{b}(\Gamma;E)\to H^{n}_{b}(\Lambda;E)$ is injective [Mon01, Proposition 8.6.2]. Therefore if $\Lambda$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic, then so is $\Gamma$. Moreover, by using Eckmann–Shapiro induction [Mon01, Proposition 10.1.3 (i)], one can show that if $\Gamma$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic, then so is $\Lambda$: Indeed, for every dual separable $\Lambda$-module $E$, we have an isomorphism $H^{n}_{b}(\Lambda;E)\cong H^{n}_{b}(\Gamma;L^{\infty}(\Gamma,E)^{\Lambda})$ for all $n\geq 0$. Now $\Lambda$-invariance implies that elements of $L^{\infty}(\Gamma,E)^{\Lambda}$ are constant on $\Lambda$-cosets of $\Gamma$, that are in finite number as $\Lambda$ has finite index in $\Gamma$. The space $L^{\infty}(\Gamma,E)^{\Lambda}$ is thus isomorphic to $E^{[\Gamma:\Lambda]}$ and so is separable if $E$ is. Then by assumption
@@
for all $n\geq 1$. However this argument does not work if $\Gamma$ is only boundedly acyclic. Therefore if $\Gamma$ is a boundedly acyclic group containing a finite index subgroup that is not boundedly acyclic, then $\Gamma$ is not $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic (explicit examples of this situation are known in degree $2$ [CFFLM23, Remark 2.4]).
Problem 4.3 (Commensurability invariance of bounded acyclicity).
@@
Problem 4.4 (c. c. versus bounded acyclicity).

### v2

@@ 2.2.4. Conjugates in commuting $\mathbb{Z}$-conjugate
The group $\Gamma$ has conjugates in commuting $\mathbb{Z}$-conjugate (abbreviated $\operatorname{c.c.\mathbb{Z}\ c.}$) if there exists a subgroup $\Lambda\leq\Gamma$ and an element $t\in\Gamma$ such that:
@@ 3.1. Correctness of Figure 1
Consider the group $\operatorname{\textup{Homeo}_{c}}(\mathbb{R}^{2})$. It has commuting $\mathbb{Z}$-conjugates [CFFLM23, Corollary 5.19], as well as commuting $\mathbb{Z}/2$-conjugates [CFFLM23, Corollary 5.11 and its proof]. It is straightforward to adapt the proof to show that it has commuting $\mathbb{Z}/n$-conjugates for every $n\geq 2$.
@@
We fix a sequence $\textbf{n}=(n_{1},n_{2},\ldots)$, where each $n_{i}\in\mathbb{Z}_{\geq 2}$. Set $\Gamma_{0}$ to be an arbitrary finitely generated non-abelian group, and define by induction $\Gamma_{i+1}\coloneqq\Gamma_{i}\wr\mathbb{Z}/n_{i+1}$. Recall that this denotes the standard restricted wreath product $\oplus_{n_{i+1}}\Gamma_{i}\rtimes\mathbb{Z}/n_{i+1}$. The standard embeddings $\Gamma_{i}\to\Gamma_{i+1}$ make this into a directed system of groups, and we denote by $\Gamma_{\mathbf{n}}$ their directed union.
@@
The group $\operatorname{\textup{Homeo}_{c}}(\mathbb{R})$ is dissipated and hence has commuting $\mathbb{Z}$-conjugates by Example 2.15 (see also Example 3.8 below). We show now that it does not have commuting $\mathbb{Z}/n$-conjugates for any $n\geq 2$.
Let $H<\operatorname{\textup{Homeo}_{c}}(\mathbb{R})$ be the copy of Thompson’s group $F$ acting on $(0,1)$ and fixing the rest of $\mathbb{R}$ pointwise. Note that $H$ is finitely generated and acts minimally on $(0,1)$. If $H$ commutes with $t^{n}$, it follows in particular that $t^{n}(0,1)=(0,1)$. Since $t$ is compactly supported it preserves the orientation, and thus $t^{n}(0)=0,t^{n}(1)=1$.  Similarly we deduce that $t(0)=0$ and $t(1)=1$. Thus ${}^{t}H$ is also supported on $(0,1)=t(0,1)$ and acts minimally on it. Let now $h\in H$ be an element whose only fixed point in $(0,1)$ is $1/2$. Because ${}^{t}H$ commutes with $h$, it must preserve its fixed point set and thus fix $1/2$. This contradicts the minimality of the action of ${}^{t}H$. It follows that $H$ and ${}^{t}H$ do not commute. ∎
@@
Now the proof of Proposition 3.5 shows that there exists a finitely generated subgroup $H_{1}<\Gamma_{1}$, such that if $[H_{1},t_{1}^{n}]=1$ for some $n\geq 2$ and some $t_{1}\in\Gamma_{1}$, then $[H_{1},{}^{t_{1}}H_{1}]\neq 1$. The assumption on $\Gamma_{2}$ implies that there exists a finitely generated subgroup $H_{2}<\Gamma_{2}$ such that for all $t_{2}\in\Gamma_{2}$ there exists $p\geq 1$ such that $[H_{2},{}^{t_{2}^{p}}H_{2}]\neq 1$.
@@ 3.2. More examples with commuting $\mathbb{Z}/2$-conjugates but not commuting $\mathbb{Z}$-conjugates
We will focus on an example of great interest in $K$-theory and homological stability: The infinite general linear group $\operatorname{GL}(\mathbb{R})$. The group $\operatorname{GL}(\mathbb{R})$ has commuting cyclic conjugates (in fact commuting $\mathbb{Z}/2$-conjugates) [CFFLM23, Corollary 5.45 and its proof]. On the other hand, we are going to show that the infinite general linear group $\operatorname{GL}(\mathbb{R})$ cannot have commuting $\mathbb{Z}$-conjugates. The proof is based on the fact that the displacing element in $\operatorname{GL}(\mathbb{R})$ has to conserve the dimension (of the matrices in the finitely generated subgroup). For this reason, a similar argument does also apply to some other groups that preserve a quantity [CFFLM23], including e.g.
@@
The fact that $\operatorname{GL}(\mathbb{R})$ has commuting $\mathbb{Z}/2$-conjugates has already been proved [CFFLM23, Corollary 5.45 and its proof]. We are left to show here that $\operatorname{GL}(\mathbb{R})$ does not have commuting $\mathbb{Z}$-conjugates.
@@ 3.3. Correctness of Figure 2
The following lemma will be useful:
@@ 3.5. Mitotic groups
Despite the similarity of the problem, adapting the argument from Proposition 3.19 to show that mitotic towers do not have commuting cyclic conjugates presents a challenge. Namely, a crucial step in the proof that $\beta(\Gamma)$ does not have commuting conjugates is the fact that the image $b^{i-1}(\Gamma)\to b^{i}(\Gamma)$ intersects trivially all edge stabilizers in the corresponding Bass–Serre tree. This is however not the case with $m^{i-1}(\Gamma)\to m^{i}(\Gamma)$, where the image is identified with an edge stabilizer, and in fact can be shown to fix infinitely many edges.
@@ 4. Problems
Due to Theorem 2.18, a natural candidate of a binate group that is not $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic is the binate tower construction, since such groups do not have commuting cyclic conjugates: In fact they do not even have commuting conjugates (Proposition 3.19).
@@
Another possible approach for constructing a boundedly acyclic group which is not $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic goes via commensurability.
Lemma 4.3.
If $\Lambda<\Gamma$ is a finite index subgroup, then $\Lambda$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic if and only if $\Gamma$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
Proof.
For every Banach $\Gamma$-module $E$, the restriction $H^{n}_{b}(\Gamma;E)\to H^{n}_{b}(\Lambda;E)$ is injective [Mon01, Proposition 8.6.2]. Therefore if $\Lambda$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic, then so is $\Gamma$.
Conversely, let $E$ be a dual separable $\Lambda$-module. Eckmann–Shapiro induction [Mon01, Proposition 10.1.3 (i)] yields a dual $\Gamma$-module $L^{\infty}(\Gamma,E)^{\Lambda}$ such that $H^{n}_{b}(\Lambda;E)\cong H^{n}_{b}(\Gamma;L^{\infty}(\Gamma,E)^{\Lambda})$ for all $n\geq 0$. Let $R$ be a set of right coset representatives, so that $\Gamma=\Lambda R$. Then for all $\lambda\in\Lambda,r\in R,f\in L^{\infty}(\Gamma,E)^{\Lambda}$ it holds $f(\lambda r)=\lambda f(r)$ [Mon01, Notation 2.4.7]. So the map
$L^{\infty}(\Gamma,E)^{\Lambda}\to E^{[\Gamma:\Lambda]}\colon f\mapsto(f(r))_{r\in R}$
is an isomorphism of Banach spaces. Since $E$ is separable, this shows that $L^{\infty}(\Gamma,E)^{\Lambda}$ is separable. So if $\Gamma$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic, then
@@
for all $n\geq 1$, which shows that $\Lambda$ is $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic. ∎
The first direction also works with trivial coefficients, since the module does not change along the way: If $\Lambda$ is boundedly acyclic, then $\Gamma$ is boundedly acyclic. However, the other direction really needs the change of coefficients, which can be witnessed by explicit examples in degree $2$ [CFFLM23, Remark 2.4]. Therefore, if $\Gamma$ is a boundedly acyclic group containing a finite index subgroup that is not boundedly acyclic, then $\Gamma$ is not $\operatorname{\mathfrak{X}^{\textup{sep}}}$-boundedly acyclic.
Problem 4.4 (Commensurability invariance of bounded acyclicity).
@@
Problem 4.5 (c. c. versus bounded acyclicity).
@@
Declarations
There is no conflict of interest to declare. There is no relevant data related to the paper.